(D^3-7DD'^2-6D'^3)z=cos(2x-3y)


1
Expert's answer
2020-07-12T15:05:50-0400

"(D^3-7DD'^2-6D'^3)z=cos(2x-3y)"

Auxillary equation is:

"m^3-7m-6=0"

"(m+2)(m+1)(m-3)=0"

"m = -1, -2, 3"

C.F.="f_1(y-x)+f_2(y-2x)+f_3(y+3x)"

P.I. "=\frac 1 {D^3-7DD'^2-6D'^3}cos(2x-3y)"

Put "D^2 = -4, D'^2 = -9"

P.I. "= \frac 1 {-4D+63D+54D'}cos(2x-3y) =" "\frac 1 {59D+54D'}cos(2x-3y)"

"=\frac {59D-54D'} {3481D^2-2916D'^2}cos(2x-3y)" "= \frac {59D-54D'} {=13924+26244}cos(2x-3y)"

"=\frac 1 {12320}(59Dcos(2x-3y)-54D'cos(2x-3y))"

"=\frac 1 {12320}(-118sin(2x-3y)-162sin(2x-3y))"

"=-\frac {280}{12320}sin(2x-3y)=-\frac{7}{176}sin(2x-3y)"

The complete solution "z=" C.F.+P.I."=f_1(y-x)+f_2(y-2x)+f_3(y+3x)" "-\frac{7}{176}sin(2x-3y)"

Answer: "z=f_1(y-x)+f_2(y-2x)+f_3(y+3x)-\frac{7}{176}sin(2x-3y)"